# Data Assetization via Resources-decoupled Federated Learning

### v1

@@ Abstract.
With the development of the digital economy, data is increasingly recognized as an essential resource for both work and life. However, due to privacy concerns, data owners tend to maximize the value of data through information flow rather than direct data transfer. Federated learning (FL) provides an effective approach to collaborative training models while preserving privacy. However, different data owners not only have variations in the quantity and quality of their data resources but also face mismatches between data and computing resources as model parameters and training data grow. These challenges hinder data owners’ willingness to participate and reduce the effectiveness of data assetization. In this work, we first identify the resource-decoupled FL environment, which includes model owners, data owners, and computing centers. We design a Tripartite Stackelberg Model and theoretically analyze the Stackelberg-Nash Equilibrium (SNE) for participants to optimize global utility. We propose the Quality-aware Dynamic Resources-decoupled FL algorithm (QD-RDFL), in which we derive and solve the optimal strategies of all parties to achieve SHE using backward induction, and a dynamic optimization mechanism is designed to improve the optimal strategy profile by evaluating the contribution of data quality from data owners to the global model during real training. Our comprehensive experiments demonstrate that our method effectively encourages the linkage of the three parties involved, maximizing global utility and data asset value.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacyccs: Information systems Information systems applications
@@ 1. Introduction
With the rapid development of the digital economy, data have unlocked the potential of machine learning (ML), driving the growth of various industries and becoming a key factor of production in the new era (Pei et al., 2023). Data assets have measurable value, can be quantified in monetary terms, and generate economic benefits for organizations, such as improving productivity, optimizing business processes, and supporting decision making (Fernandez et al., 2020). Nevertheless, transforming data into assets through market transactions is clumsy. Although there are many ingenious data pricing models (Badewitz et al., 2022; Hao et al., 2024; Pei, 2022), direct data market transactions are a cumbersome way of data assetization. First, for privacy protection reasons, data owners (organizations or institutions) may not be willing to participate in data transactions and are subject to legal and regulatory restrictions, such as the General Data Protection Regulation (GDPR) (GDPR, 2018), which requires that data remain where it is generated. Secondly, data owners are more concerned about how the information (knowledge) embedded in the data can improve their business efficiency (Bi et al., 2024). Therefore, they are more inclined to enable the circulation of information to maximize the value of the data rather than direct data transactions. In recent years, federated learning (FL) has emerged as a successful paradigm for collaborative model training under privacy constraints (McMahan et al., 2017; Yang et al., 2019b). It is widely adopted and benefits the finance, insurance, and healthcare industries (Yang et al., 2019a). FL maintains local data while training a global model collaboratively, avoiding direct data transfer. It offers an effective way to data assetization through the circulation of information.
@@
However, the inherent challenge to maximizing data assets’ potential value through FL training high-quality global models is heterogeneous data resources, i.e., data owners exhibit differences in data quantity and quality (Luo et al., 2024; Shi et al., 2024a; Sun et al., 2024). Since these features of the data are key factors in determining the effectiveness of ML models, this disparity may cause high-quality data owners to lack interest in participating in FL. Moreover, a more pressing challenge is the mismatch between data and computing resources. With recent advances in artificial intelligence, the positive correlation between model performance and the number of model parameters has been well-established (Kaplan et al., 2020; Brown et al., 2020; Devlin et al., 2019). Recently, Large models have garnered significant attention in academia and industry because of their outstanding performance. However, as the number of parameters increases, data owners face the dilemma of insufficient computing resources. We conducted preliminary experiments with different computing resource configurations (the model size is 3B, 7B, and 13B) and training durations required for a fixed amount of data (140B tokens) according to (Kaplan et al., 2020). The results shown in Tab. 1 indicate that within acceptable training times, the lower limit of client-side computing resources exceeds most data owners’ typical computing capabilities. We refer to this real-world scenario as resource-decoupled resources FL, where data resources with unbalanced quality and quantity are distributed across different data owners and a mismatch between data resources and computing resources, as shown in Fig. 1.
@@
Table 1. GPU Configurations, Model Size, and Training Performance
@@
Focusing on the problem of heterogeneous data resources in FL, researchers have developed advanced incentive mechanisms that employ game theory, contracts, and auctions for interactions between cloud servers and multiple clients (Hu et al., 2024; Zhan et al., 2020; Huang et al., 2024; Zeng et al., 2020; Zheng et al., 2024; Sun et al., 2022). Among them, game theory formulates the interaction between incentive structures and uses mathematical theories to solve individual optimization strategies in complex environments with competitive phenomena (such as FL). Stackelberg games provide solutions for non-cooperative games and are applied to FL scenarios with multiple actors (Bi et al., 2024; Luo et al., 2024; Hu et al., 2024). There are also many dynamic game solutions for complex decision processes (Tang et al., 2024; Lin et al., 2024). Many researchers measure the contribution of client data and provide a fair solution for profit distribution, incentivizing data owners to participate in training and improving the performance of the final model. Some of them use metrics such as validation accuracy and loss values derived from training processes (Hu et al., 2024; Baghcheband et al., 2024; Sun et al., 2024). Moreover, the Shapley values successfully quantify the marginal contribution of data in ML (Tarun et al., 2024; Shi et al., 2024b; Hu et al., 2024) and provide a fair incentive solution for cooperative games. However, the classical Shapley value calculation is computationally expensive and impractical in FL. To address this, efficient approximations of marginal contribution calculation have been proposed (Yang et al., 2024). Although these game-theoretic incentive mechanisms have achieved results in server-client FL, they cannot be directly applied to resource-decoupled FL scenarios.
@@
Unlike previous server-client models, we identify three roles in the resource-decoupled FL: model owner (coordinator), data owners, and computing centers. In Fig. 1, the model owner initiates FL training to build an optimal global model that maximizes the overall utility. Data owners hold data with varying quality and quantity. Their objective is to transform their data into advanced knowledge to create business value, such as models or monetary profits, ultimately maximizing the personal data asset. computing centers are large-scale data centers with abundant computing resources seeking to increase their revenue by leasing them. For a specific business scenario, a model owner coordinates $N$ data owners and $M$ computing centers in FL. To incentivize three-party collaboration, a reward scheme is designed. This scheme maximizes the utility of the global model, the profit of personal data, and the rental income of computing resources. It encourages data owners to contribute high-quality and high-quantity data while computing centers align their computing resources with the data volume to achieve maximum revenue. Our goal is to maximize the value of data assets. According to the above analysis, when the utility of the global model reaches its maximum, all parties can achieve higher profits, ensuring optimal global utility. Therefore, the new challenges to face include: (1) How can FL be effectively implemented in scenarios where data owners and computing centers are separated? (2) How can an incentive mechanism be designed that encourages all parties to adopt optimal strategies, thus achieving global utility optimization? (3) How do we assess the contribution of data owners to the final model and quantify their rewards?
@@
Motivated by this, we focus on the data assetization approach, in which data do not flow, but the information does. It constructs a resource-decoupled FL framework and designs a Tripartite Stackelberg Model to encourage three parties to optimize their profit, in which the model owner acts as the leader, the data owners as the sub-leaders, and the computing centers as the followers. We derive three utility functions according to each one’s profit and analyze the Nash equilibrium of the sub-game models. Global utility optimization is defined as achieving the Stackelberg-Nash Equilibrium (SNE), and backward induction is introduced to solve the model, resulting in three-party strategies that maximize global utility. We evaluate the model contribution of data owners according to data quality, and an adjustment mechanism is designed to improve the optimal strategy, encouraging data owners to provide high-quality and quantity data in FL training to optimize global model performance.
Specifically, our contributions are as follows:
@@
  • We propose a framework for resource-decoupled FL involving model owners, data owners, and computing centers. We define utility functions by linking data value, model value, and monetary value to optimize global utility, thus maximizing data asset value.
@@
  • We develop a Tripartite Stackelberg model and analyze SNE of the model. Using backward induction, we derive optimal strategies for participants to maximize global utility.
@@
  • We introduce the Quality-aware Dynamic Resource-decoupled FL algorithm (QD-RDFL), in which a dynamic optimization mechanism is designed to improve the optimal strategies by evaluating the contribution of data owners.
@@
  • We conduct extensive experiments that demonstrate the proposed method effectively encourages participants, maximizing global utility and data asset value.
@@ 2. Related Works
Data Valuation in ML: To encourage participants to contribute higher quality data and actively participate in FL, accurately assessing the contributions of participants is critical for an equitable distribution of benefits. Existing research primarily evaluates contributions based on the performance of the final model (Zhan et al., 2020; Huang et al., 2024; Zeng et al., 2020). Some studies rely on previous data provided by clients or subjectively define data quality (Luo et al., 2024; Shi et al., 2024b; Zheng et al., 2024; Sun et al., 2022). In contrast, others assess contributions through actual performance metrics during training, such as validation accuracy, loss values, and Shapley values (Hu et al., 2024; Baghcheband et al., 2024; Tarun et al., 2024). To objectively quantify the contributions of data owners in FL and design reward mechanisms that encourage the provision of high-quality data, we evaluate the data value based on actual training performance, considering data quantity and quality.
Notations Description
@@
$d_{m}$ Data quantity undertaken by $C_{m}$
@@
Table 2. Notations and Descriptions
Incentive Mechanisms in FL: Game-theoretical incentive mechanisms simulate strategic interactions between the server and the clients and the competitive dynamics among clients. As a result, these mechanisms have garnered extensive attention in FL. The classical Shapley value provides a robust framework to assess each participant’s contribution, which can be leveraged to design incentive mechanisms for cooperative games (Shi et al., 2024a; Yang et al., 2024). However, traditional Shapley values are computationally intensive, increasing the workload of FL training. Incentive mechanisms based on Stackelberg games position the server as the leader and clients as followers, typically maximizing the utility for both parties. For example, (Hu et al., 2024) designed a three-layer Stackelberg game for FL involving cloud-edge servers and clients, optimizing social utility through Pareto utility enhancements. (Huang et al., 2024; Zheng et al., 2024) used a two-stage Stackelberg game-based incentive mechanism, where the server first engages in a Stackelberg game to maximize its utility, followed by clients participating in a non-cooperative game to achieve their utility maximization through Nash equilibrium solutions. Additionally, (Tang et al., 2024) introduced a dynamic game solution for complex decision-making processes between servers and devices in the FL process as a continuous iterative game to optimize social welfare. (Lin et al., 2024) constructed a two-layer game architecture, where the upper layer mechanism dynamically adjusts selection strategies through reinforcement learning to achieve long-term utility maximization, and the lower layer mechanism benefits individuals through evolutionary games.
3. Methodology
3.1. Problem Statement
This work focuses on maximizing the data assets’ value through FL in a resource-decoupled scenario. The ultimate profit of the three parties is linked to the quality of the global model. However, predicting the model’s quality is challenging. Therefore, inspired by (Hu et al., 2024), we define the problem as a global utility optimization in resource-decoupled FL.In this section, we first formalize the global optimization problem and then provide the utility functions of three parties. The important notations are listed in Tab. 2.
Global Utility Optimization: We will design an incentive mechanism to derive the optimization strategy for the three parties, with the aim of optimizing global utility $U_{Global}$. The sub-optimization problems include the utility $U_{s}$ of model owner $S$, the utility $U_{n}$ of data owner $D_{n}$, and the utility $U_{m}$ of computing center $C_{m}$. As in the previous analysis, the overall utility $\sum_{N}U_{n}$ of data owners and the overall utility $\sum_{M}U_{m}$ of the computing centers increase when an optimum $U_{s}$ is obtained. Since $U_{s}$ is up to the quality of the global model, we need to evaluate the contribution of quantity and quality of data provided by data owners to global model updates. These evaluations encourage data owners to provide higher quality model updates, thus improving the global model. However, due to the separation of computing and data resources, FL tasks require data owners to be matched with computing centers. Consequently, computing centers maximize their utility by optimally matching with the data owners.
In resource-decoupled FL, the computing center acts as the ’client’ in classical FL, such as FedAvg (Yang et al., 2019b), after matching with the specific data owner. Therefore, the resource-decoupled FL aims to train an optimal global model $w^{*}$, minimizing the global loss function $F(w)$ by aggregating the high quality local update $F_{i}(w)$, as shown in Equ. 1.
@@ 3.1. Problem Statement
where $K$ represents the number of computing centers matching successful. We assume $K\leq N\leq M$ in this work. $p_{i}=\frac{x_{n}}{\sum_{K}x_{n}}$ is data provided ratio of data owner $D_{n}$.
To encourage the three parties, we analyze and formalize the utility of each computing center, data owner, and model owner as the basis of the reward distribution. We evaluate the model contribution of data owner based on the data quality and quantity and transform the data value into model value. Meanwhile, considering both cost and profits, we establish the association between data value, model value, and monetary value and derive the utility functions of the three parties.
Utility of Computing Center: The utility of computing centers is the difference between the cost and profit gained from renting out computing power. Since the profit is related to the matching data quantity provided by the data owner, the larger the quantity, the higher the profit. However, the cost of completing the same task varies due to differences in hardware devices and computing power across computing centers. The goal of $C_{m}$ is to determine the optimal matching with a specific data owner according to undertaking data quantity $d_{m}$ to maximize its utility $U_{m}$. Accordingly, we define the utility of the computing center $C_{m}$ as Equ. 3.
@@
where $\sigma_{m}$ is the computational power factor of $C_{m}$, which reflects the difference in cost across various computing centers. $\epsilon$ is the unit training cost of the computing center. $\rho$ is the unit training payment of the data owner. $\lambda$ is the market regulating factor to adjust the two-sided marketplace between data owners and computing centers. When $M<N$, $\lambda$ is a constant greater than 1; when $M>N$, $\lambda$ is less than 1; and when $M=N$, $\lambda$ is 1.
Utility of Data Owner: The utility of data owner is the difference of the model profit generated by their data with the cost of model training. Specifically, the model profit of personal data originates from the total model payment provided by model owners, while the cost arises from the fee for renting computing center per training unit. Followed by (Sun et al., 2024; Hu et al., 2024), we evaluate the contribution $q_{n}$ of each data owner to the global model based on the quantity $x_{n}$ and quality $f_{n}$ of their training data. Thus, given a certain total model payment $\eta$, the higher the contribution, the greater the reward data owners receive. This encourages data owners to provide higher quality and larger quantity of data in FL. The utility $U_{n}$ of data owner $D_{n}$ is defined as shown in Equ. 5.
@@
Utility of Model Owner: The utility of the model owner is the difference between global model profit and the total model payment $\eta$ to data owners. As known, the higher the model contributions from the data owners, the higher the quality of the global model, i.e., global model profit. Therefore, we define the quality of the global model as a concave function $\alpha\cdot g(\sum^{N}_{n=1}q_{n})$ which depends on the model contributions of all data owners, where $\alpha>0$ is a preset parameter. The utility of model owner $S$ is defined as shown in Equ. 7.

### v2

@@ Abstract.
With the development of the digital economy, data is increasingly recognized as an essential resource for both work and life. However, due to privacy concerns, data owners tend to maximize the value of data through the circulation of information rather than direct data transfer. Federated learning (FL) provides an effective approach to collaborative training models while preserving privacy. However, as model parameters and training data grow, there are not only real differences in data resources between different data owners, but also mismatches between data and computing resources. These challenges lead to inadequate collaboration among data owners, compute centers, and model owners, reducing the global utility of the three parties and the effectiveness of data assetization. In this work, we first propose a framework for resource-decoupled FL involving three parties. Then, we design a Tripartite Stackelberg Model and theoretically analyze the Stackelberg-Nash equilibrium (SNE) for participants to optimize global utility. Next, we propose the Quality-aware Dynamic Resources-decoupled FL algorithm (QD-RDFL), in which we derive and solve the optimal strategies of all parties to achieve SNE using backward induction. We also design a dynamic optimization mechanism to improve the optimal strategy profile by evaluating the contribution of data quality from data owners to the global model during real training. Finally, our extensive experiments demonstrate that our method effectively encourages the linkage of the three parties involved, maximizing the global utility and value of data assets.
@@
ccs: Security and privacyccs: Information systems Information systems applications
@@ 1. Introduction
With the rapid development of the digital economy, data have unlocked the potential of machine learning (ML), driving the growth of various industries and becoming a key factor of production in the new era (Pei et al., 2023). Data assets have measurable value, can be quantified in monetary terms, and generate economic benefits for organizations, such as improving productivity and optimizing business processes (Fernandez et al., 2020). However, transforming data into assets through market transactions is clumsy. Although there are many ingenious data pricing models (Badewitz et al., 2022; Hao et al., 2024; Pei, 2022), direct data market transactions are a cumbersome way of data assetization. First, for privacy protection reasons, data owners (organizations or institutions) may not be willing to participate in data transactions and are subject to legal and regulatory restrictions, such as the General Data Protection Regulation (GDPR) (GDPR, 2018), which requires that data remain where it is generated. Secondly, data owners are more concerned about how the information (knowledge) embedded in the data can improve their business efficiency (Bi et al., 2024). Therefore, they are more inclined to enable the circulation of information to maximize the value of the data rather than direct data transactions. In recent years, federated learning (FL) has emerged as a successful paradigm for collaborative model training under privacy constraints (McMahan et al., 2017; Yang et al., 2019a). It is widely adopted and benefits the finance, insurance, and healthcare industries (Yang et al., 2019b; Zhang et al., 2021; Wang and Feng, 2024; Rieke et al., 2020). FL keeps data locally while training a global model collaboratively, thus offering an effective way for data assetization.
@@
However, to maximize the potential value of data assets by FL training high-quality global models, the inherent challenge is heterogeneous client data resources, i.e., data owners exhibit differences in data quantity and quality (Luo et al., 2024; Shi et al., 2024; Sun et al., 2024). Since quantity and quality of data are key factors in determining the effectiveness of ML models, the data heterogeneity can cause high-quality data owners to lose the motivation to participate in FL. Moreover, a more pressing challenge is the mismatch between data and computing resources. With recent advances in artificial intelligence, the positive correlation between model performance and the number of model parameters has been well-established (Brown et al., 2020; Devlin et al., 2019). Large language models have recently garnered significant attention in academia and industry because of their outstanding performance. However, as the number of parameters increases, data owners face the dilemma of insufficient computing resources. We conducted preliminary experiments for different computing resource configurations, model sizes (3B, 7B, and 13B), and training durations required with a fixed amount of data (140B tokens) (Kaplan et al., 2020). The results in Tab. 1 indicate that within acceptable training times, the lower limit of computing resources exceeds most data owners’ typical computing capabilities. We refer to this real-world scenario as resource-decoupled FL, where data resources with unbalanced quality and quantity are distributed across different data owners and a mismatch between data and computing resources, as shown in Fig. 1.
@@
Table 1. GPU Configurations, Model Size, and Training Performance
@@
Focusing on the problem of heterogeneous data resources in FL, researchers have developed advanced incentive mechanisms that employ game theory, contracts, and auctions for interactions between the cloud server and multiple clients (Hu et al., 2024; Zhan et al., 2020; Huang et al., 2024; Zeng et al., 2020; Zheng et al., 2024; Sun et al., 2022). Among them, game theory formulates the interaction between incentive structures and uses mathematical theories to solve individual optimization strategies in complex environments with competitive phenomena (such as FL). Stackelberg games provide solutions for non-cooperative games and are applied to FL scenarios with multiple roles (Bi et al., 2024; Luo et al., 2024; Hu et al., 2024). There are also many dynamic game solutions for complex decision processes (Tang et al., 2024; Lin et al., 2024). Many researchers measure the contribution of client data and provide a fair solution for profit distribution, encouraging data owners to participate in FL and improving the performance of the global model. Some metrics are used to evaluate the contribution, such as verification accuracy and loss values (Hu et al., 2024; Baghcheband et al., 2024; Sun et al., 2024). Moreover, the Shapley values successfully quantify marginal contributions (Tarun et al., 2024; Shi et al., 2024; Hu et al., 2024) and provide a fair incentive solution for cooperative games. However, the classical Shapley value calculation is computationally expensive and impractical in FL. Therefore, efficient approximations Shapley calculation have been proposed (Yang et al., 2024). Although these game-theoretic incentive mechanisms have achieved results in server-client FL, they cannot be directly applied to resource-decoupled FL scenarios. Due to the separation of computing and data resources, the game should be complicated by introducing new roles.
@@
Unlike previous server-client models, we identify three roles in the resource-decoupled FL: model owner (coordinator), data owners, and computing centers. In Fig. 1, the model owner initiates FL training to build an optimal global model that maximizes the overall utility. Data owners hold data of varying quality and quantity. Their objective is to transform their data into advanced knowledge to create business value, such as models and monetary benefits, ultimately maximizing personal data assets. Computing centers are large-scale data centers with abundant computing resources that seek to increase their rental income. For a specific business scenario, the model owner coordinates $N$ data owners and $M$ computing centers in FL. In this work, we are dedicated to maximizing the value of the data assets. To incentivize tripartite collaboration, a reward scheme should be designed. This scheme encourages data owners to contribute high-quality and high-quantity data while computing centers align their computing resources with the data volume to achieve maximum revenue. Therefore, it maximizes the profit of personal data, the rental income of computing resources, and the utility of the global model. In light of the analyses above, when the utility of the global model reaches its maximum, all parties can achieve higher profits, ensuring optimal global utility. Therefore, the new challenges include: (C1) How can FL be effectively implemented in scenarios where data owners and computing centers are separated? (C2) How can an incentive mechanism be designed that encourages all parties to adopt optimal strategies, thus achieving global utility optimization? (C3) How do we assess the contribution of data owners to the final model and quantify their rewards?
@@
Motivated by this, we focus on the approach of data assetization, that data do not circulate but the information does. We first propose a resource-decoupled FL framework for solving (C1), in which the model owner coordinates the FL training of the three parties, the data centers are in charge of providing the data, and the matching data centers are responsible for local training. Then, we design a Tripartite Stackelberg Model to encourage three parties to optimize their profit, thus solving (C2), in which the model owner is the leader, the data owners are sub-leaders, and the computing centers are followers. We derive three utility functions according to each one’s profit and analyze the Nash equilibrium of the sub-game models. Global utility optimization is defined as the achievement of the Stackelberg-Nash equilibrium (SNE), and backward induction is introduced to solve the model, resulting in three-party strategies that maximize global utility. Finally, to overcome (C3), we evaluate the model contribution of data owners according to data quality, and an adjustment mechanism is designed to improve the optimal strategy, encouraging data owners to provide high-quality and quantity data in FL training to optimize the global model’s performance.
Specifically, our contributions are as follows.
@@
  • We propose a resource-decoupled FL framework that involves model owners, data owners, and computing centers and define utility functions and the global utility optimization problem to maximize the value of data assets.
@@
  • We develop a Tripartite Stackelberg model and analyze the SNE of the model. Using backward induction, we derive optimal strategies to maximize global utility.
@@
  • We introduce the Quality-aware dynamic resource-decoupled FL algorithm (QD-RDFL), in which a dynamic optimization mechanism is designed to improve optimal strategies by evaluating the contribution of data owners.
@@
  • The comprehensive experiments demonstrate that our method effectively encourages participants, maximizing global utility and the value of data assets.
@@ 2. Related Works
Data Valuation in ML: To encourage participants to contribute higher quality data and actively participate in FL, accurately assessing the contributions of participants is critical for an equitable distribution of benefits. Existing research primarily evaluates contributions based on the performance of the final model (Zhan et al., 2020; Huang et al., 2024; Zeng et al., 2020). Some studies rely on previous data provided by clients or subjectively define data quality (Luo et al., 2024; Shi et al., 2024; Zheng et al., 2024; Sun et al., 2022). In contrast, others assess contributions through actual performance metrics during training, such as validation accuracy, loss values, and Shapley values (Hu et al., 2024; Baghcheband et al., 2024; Tarun et al., 2024). To objectively quantify the contributions of data owners in FL and design reward mechanisms that encourage the provision of high-quality data, we evaluate the data value based on actual training performance, considering data quantity and quality.
Incentive Mechanisms in FL: Game-theoretical incentive mechanisms simulate strategic interactions between the server and the clients and the competitive dynamics among clients. As a result, these mechanisms have garnered extensive attention in FL. The classical Shapley value provides a robust framework to assess each participant’s contribution, which can be leveraged to design incentive mechanisms for cooperative games (Shi et al., 2024; Yang et al., 2024). However, traditional Shapley values are computationally intensive, increasing the workload of FL training. Incentive mechanisms based on Stackelberg games position the server as the leader and clients as followers, typically maximizing the utility for both parties. For example, (Hu et al., 2024) designed a three-layer Stackelberg game for FL involving cloud-edge servers and clients, optimizing social utility through Pareto utility enhancements. (Huang et al., 2024; Zheng et al., 2024) used a two-stage Stackelberg game-based incentive mechanism, where the server first engages in a Stackelberg game to maximize its utility, followed by clients participating in a non-cooperative game to achieve their utility maximization through Nash equilibrium solutions. Additionally, (Tang et al., 2024) introduced a dynamic game solution for complex decision-making processes between servers and devices in the FL process as a continuous iterative game to optimize social welfare. (Lin et al., 2024) constructed a two-layer game architecture, where the upper layer mechanism dynamically adjusts selection strategies through reinforcement learning to achieve long-term utility maximization, and the lower layer mechanism benefits individuals through evolutionary games.
3. Methodology
3.1. Problem Statement
This work focuses on maximizing the value of data assets through FL in a resource-decoupled scenario. The ultimate profit of the three parties is linked to the quality of the global model. However, predicting the model’s quality is challenging. Therefore, we define the problem as a global utility optimization in resource-decoupled FL.In this section, we first formalize the global optimization problem and then provide the utility functions of three parties. The necessary notations are listed in Tab. 2.
Table 2. Notations and Descriptions
Notations Descriptions
@@ 3.1. Problem Statement
$d_{m}$ , $\left|d_{m}\right|$ Data quantity and max quantity undertaken by $C_{m}$
@@
$L$ Optimal adjustment round
Global Utility Optimization: We will design an incentive mechanism to derive the optimization strategy for the three parties to optimize global utility $U_{Global}$. The sub-optimization problems include the model owner’s utility $U_{s}$, the utility $U_{n}$ of data owner $D_{n}$, and the utility $U_{m}$ of computing center $C_{m}$. As in the previous analysis, the overall utility $\sum_{N}U_{n}$ of data owners and the overall utility $\sum_{M}U_{m}$ of the computing centers increase when an optimum $U_{s}$ is obtained. Since $U_{s}$ is up to the quality of the global model, we need to evaluate the contribution of quantity and quality of data provided by data owners to global model updates. These evaluations encourage data owners to provide higher-quality model updates, thus improving the global model. However, due to the separation of computing and data resources, FL tasks require data owners to be matched with computing centers. Consequently, computing centers maximize their utility by optimally matching with the data owners.
In resource-decoupled FL, the computing center acts as the ’client’ in classical FL, such as FedAvg (Yang et al., 2019a), after matching with the specific data owner. Therefore, the resource-decoupled FL aims to train an optimal global model $w^{*}$, minimizing the global loss function $F(w)$ by aggregating the high-quality local update $F_{i}(w)$, as shown in Equ. 1.
@@
where $K$ represents the number of computing centers matching successfully. We assume $K\leq N\leq M$ in this work. $p_{i}=\frac{x_{n}}{\sum_{K}x_{n}}$ is data provided ratio of data owner $D_{n}$.
To encourage the three parties, we analyze and formalize the utility of each computing center, data owner, and model owner as the basis of the reward distribution. We evaluate the model contribution of the data owner based on the data quality and quantity and transform the data value into model value. Meanwhile, considering both cost and profits, we derive the utility functions of the three parties.
Utility of Computing Center: The utility of computing center $U_{m}$ is the difference between the cost and profit gained from renting out computing power. Since the profit is related to the matching data quantity provided by the data owner, the larger the quantity, the higher the profit. However, the cost of completing the same task varies due to differences in hardware devices and computing power across computing centers(Strubell et al., 2019; Sze et al., 2017; Samsi et al., 2023). The goal of $C_{m}$ is to determine the optimal matching with a specific data owner according to undertaking data quantity $d_{m}$ to maximize its utility $U_{m}$. Accordingly, we define $U_{m}$ of $C_{m}$ as Equ. 3.
@@
where $\sigma_{m}$ is the computational power factor of $C_{m}$, which reflects the difference in cost across various computing centers. $\epsilon$ is the unit training cost of the computing center. $\rho$ is the unit training payment of the data owner. $\lambda$ is the market regulating factor to adjust the two-sided marketplace between data owners and computing centers. In this work, we assume $M=N$; at this point, $\lambda$ is 1.
Utility of Data Owner: The utility of data owner $U_{n}$ is the difference between the model profit generated by their data and the cost of model training. Specifically, the model profit of personal data originates from the total model payment provided by model owners, and the cost arises from the fee for renting a computing center per training unit. Following (Sun et al., 2024; Hu et al., 2024), we evaluate the contribution $q_{n}$ of each data owner to the global model based on the quantity $x_{n}$ and quality $f_{n}$ of their training data. Thus, given a certain total model payment $\eta$, the higher the contribution, the higher the reward data owners receive. The evaluations encourage data owners to provide higher quality and larger quantity of data in FL. The $U_{n}$ of $D_{n}$ is defined as shown in Equ. 5.
@@
Utility of Model Owner: The utility of the model owner $U_{s}$ is the difference between the global model profit and the total model payment $\eta$ to data owners. As known, the higher the model contributions from the data owners, the higher quality of the global model, i.e., global model profit. Therefore, we define the quality of the global model as a concave function $\alpha\cdot g(\sum^{N}_{n=1}q_{n})$ which depends on the model contributions of all data owners, where $\alpha>0$ is a preset parameter. The $U_{s}$ of model owner $S$ is defined as Equ. 7.